OpenStudy (thesecret20111):

Is the point (1,4) defined by y<2x+1?

OpenStudy (callisto):

Consider LHS y = 4 Consider RHS 2x+1 = 2(1)+1 = 3 < LHS So it's not.. I guess :P

OpenStudy (anonymous):

By convention, when we write a pair of numbers as (a,b), we mean that the first one is a fixed x and the second one is a fixed y. Moreover, y<2x+1 geometrically means all points under the line y=2x+1. You plug in 1 for x and 4 for y, then 4>2.1+1, meaning (1,4) is over the line y=2x+1. Therefore, this point is not defined by the given inequality.

OpenStudy (thesecret20111):

Thanks so much for the fast replys it really helps! This sites awesome :D